OpenStudy (anonymous):

write y=-3x^2+12x-21 in vertex form A)y=-3(x-2)^2-9 B)y=-3(x+2)^2-9 C)y=-3(x+2)^2+9 D)y=-3(x-2)^2+9

OpenStudy (anonymous):

Do you know the basic vertex form? That is without the integers in the equation?

OpenStudy (anonymous):

no i do not, sorry

OpenStudy (anonymous):

f(x) = a(x-h)^2 + k this is vertex form. But to complete it, you have to complete the square. Do you remember that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Okay, so right now you have your equation in general form. To complete the square you take the x integer and divide by 2. After you divide by two, you square the result to recieve the "completed square"

OpenStudy (anonymous):

So take 12/2 = 6. 6^2 = 36

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

One second, looking at my notes. Its been a while since I have found the vertex myself!

OpenStudy (anonymous):

Okay now that u have that you have take subtract the 36 from the other side to keep things equal. Let me solve this so I know what you get

OpenStudy (anonymous):

36 from 21?

OpenStudy (anonymous):

You must add the 21 to zero and subtract 36 from the zero. So it will be -3x^2 +12x = 21 - 36

OpenStudy (anonymous):

When you combine those two to get -15. Now you can pull the 3 out. -3(x-2)^2 - 9

OpenStudy (anonymous):

where do you get -9 from?